# ToolACE: Winning the Points of LLM Function Calling

### v1

@@ Abstract
Function calling significantly extends the application boundary of large language models, where high-quality and diverse training data is critical for unlocking this capability. However, real function-calling data is quite challenging to collect and annotate, while synthetic data generated by existing pipelines tends to lack coverage and accuracy. In this paper, we present ToolACE, an automatic agentic pipeline designed to generate accurate, complex, and diverse tool-learning data. ToolACE leverages a novel self-evolution synthesis process to curate a comprehensive API pool of 26,507 diverse APIs. Dialogs are further generated through the interplay among multiple agents, guided by a formalized thinking process. To ensure data accuracy, we implement a dual-layer verification system combining rule-based and model-based checks. We demonstrate that models trained on our synthesized data, even with only 8B parameters, achieve state-of-the-art performance on the Berkeley Function-Calling Leaderboard, rivaling the latest GPT-4 models. Our model and a subset of the data are publicly available at https://huggingface.co/Team-ACE/.
@@ 1 Introduction
Equipping Large Language Models (LLMs) with external tools has significantly enhanced the capability of AI Agents to solve complex real-world tasks [10, 15, 16]. The integration of function calling enables LLMs to access up-to-date information, perform delicate computations, and utilize third-party services, thereby unlocking a wide range of potential applications across various fields, e.g., workflow automation [26], financial reporting [20], and travel planning [6].
Driven by practical applications, function calls can be quite complicated due to the diversity, complexity, and interactions of real-world APIs [15]. [^1] In this report, APIs, tools, functions, and plugins are interchangeable. For instance, real-world API parameters often extend beyond simple strings or numbers to include complex types like lists, dictionaries, nested parameters, and even combinations of these. The number of these API parameters can range from zero to dozens, and the application domains of these APIs span a wide range of businesses and industries [24]. Moreover, a single API is often insufficient to complete a task; instead, multiple tools need to be used together to perform real-world tasks [10]. The input of one API may even depend on the output of another [15], making function calls even more complicated and challenging.
Considering this diversity and complexity, function calls for specific tasks or user queries can generally be grouped into three categories: single function calls, parallel function calls, and dependent function calls. In a single function call, the LLM selects and executes one function to accomplish the user’s task. For parallel function calls, the LLM performs multiple independent function calls simultaneously within one turn. Whereas dependent function calls involve the LLM making a series of sequential calls, with each subsequent call relying on the output of the previous ones.
However, current tool-augmented LLMs primarily focus on simple function-calling scenarios with limited diversity and complexity [16]. Table 1 provides an overview of the data statistics utilized in these representative tool-augmented LLMs. Although satisfying performance has been achieved on the single function calling that executes one API in one turn, the capabilities of parallel and dependent function calls are largely overlooked. Furthermore, the constrained API domains, simplistic parameter types, and uniform data formats may hinder the applicability of the function calling to more complex, real-world tasks. Executing real-world function calls requires precise API selection and accurate parameter configuration, both of which are closely tied to the accuracy of the underlying data. Yet ensuring data accuracy for tooling remains a challenging problem, particularly when dealing with diverse and complex tasks.
@@
Gorilla [14] 1645 3
ToolAlpaca [19] 3938 50
ToolLLM [15] 16464 49
Functionary [12] n/a n/a
xLAM [11] 3673 21
Granite [1] n/a n/a
@@
Therefore, in this report, we present ToolACE, a systematic tool-learning pipeline that encompasses both synthetic data generation and data verification, demonstrating robust performance with enhanced accuracy, diversity, and complexity.
Diversity. Exposing LLMs to diverse function-calling scenarios facilitates a more well-rounded cognitive skill set of tool usage [25]. In ToolACE, we propose to emphasize three dimensions of diversity for function calling: tool diversity, format diversity, and dialog diversity. Tool diversity is achieved through our tool self-evolution synthesis method (TSS), which synthesizes tools from various domains with diverse data types and constraints. For format diversity, we propose a tool format generalization method to support any mainstream tool description and tool calling format (e.g., Json, YAML, XML, markdown) in ToolACE. Dialog diversity includes simple, parallel, dependent function calls, and non-tool-use dialogs, encompassing a wide range of function-calling use cases.
Complexity. In ToolACE, we emphasize the complexity of the data, where the instruction-following data should possess sufficient complexity to develop the necessary skills for function calling. We also find that the LLMs learn more effectively when the data complexity is slightly above their current capability level. Data that is either too simple or too complex proves unproductive for LLMs. This aligns with the Zone of Proximal Development (ZPD) theory in educational psychology, which states that learning is most effective when the tasks are just beyond the learner’s current level of independent capability but achievable with suitable support [17].
Accuracy. Data accuracy is fundamental to the effectiveness of tool-augmented LLMs. To achieve high-quality data, we first implement a formalized thinking and self-consistency strategy to enhance reliability during data generation. Then we deploy a dual-layer verification (DLV) system, integrating rule-based and model-based checks, to further improve accuracy.
@@ 2 Data Generation Pipeline
Effective use of synthetic data significantly enhances the capabilities of large language models (LLMs) [13]. Hence, in ToolACE, we propose an automated agentic framework for tool learning to generate high-quality, diverse, and complex data using advanced LLMs, as illustrated in Figure 1. The proposed framework deploys various agents to recursively synthesize diverse APIs, collaboratively construct formalized dialogs, and rigorously reflect on data quality. The following sections present our Tool Self-evolution Synthesis (TSS) module, Multi-Agent Interactive Dialog Generation (MAI) module, and Dual-Layer Validation Process (DLV).
@@ 2.1 Tool Self-evolution Synthesis
The variety of APIs significantly underpins the diversity and complexity of function-calling data. As shown in Table 1, ToolACE has established a comprehensive API pool that surpasses other representative tool-augmented LLMs in both quantity and domain coverage, incorporating both real and synthesized APIs. Beyond collecting real API data, we developed a Tool Self-Evolution Synthesis (TSS) module that synthesizes API definitions with various data types and constraints. Specifically, we utilize pretraining data to extract an API context tree, where each node represents a potential application domain and functionalities for function calls, e.g., finance, health, and transport. Combining the sampled functionalities and an API example, an agent powered by a frontier LLM synthesizes new APIs. The diversity and complexity of the APIs are gradually increased through recursive self-evolution and updates. This process is achieved through three major steps: 1) Speciation, 2) Adaption, and 3) Evolution. The detailed process is illustrated in Figure 2.
@@
Figure 1: The overall framework of ToolACE, which mainly consists of Tool Self-evolution Synthesis (TSS), Multi-Agent Interactive Dialog Generation (MAI), and Dual-Layer Validation Process (DLV).
Speciation.
@@
APIs with extensive domain coverage enable tool-augmented LLMs to learn a wider array of use cases from various applications and industries, thereby significantly enhancing their generalization ability. In the speciation step, we propose to create a hierarchical API context tree to guide the synthesis process with possible API domains and functionalities.
@@
We observe that the pretraining data for LLMs encompasses one of the most diverse sources of human corpus, providing a solid foundation for extracting various API domains and use cases. Starting with API-related raw documents from the pretraining data (e.g., technical manuals, API documentation, product specifications, user guides, and tutorials), we prompt an LLM to extract an API domain along with all possible API functionalities or use cases from each document. Children nodes of the context tree are recursively generated at each step. Our final context tree encompasses a comprehensive structure that covers 30 primary domains (e.g., Entertainment, Education), divided into 390 coarse-grained sub-domains (e.g., Music, Anime, Books), and further refined into 3,398 distinct fine-grained domains (e.g., Music Streaming, Live Music). This extensive context tree includes an estimated hundred thousand specific functionalities. The left part of Figure 2 showcases the subtree under the entertainment domain as an example.
Adaption.
@@
In the adaption step, we specify the domain and complexity level of each API. To determine the complexity level of individual APIs, we sample a subtree from the child nodes at the fine-grained domain level of the API context tree and obtain unique functionalities, ensuring that APIs within the same domain possess distinct functionalities. Specifically, a more complex API should cover more context tree nodes, thereby acquiring more domain-specific and detailed capabilities. Conversely, a simpler API may include only a single sub-node from the context tree, focusing on an easy, straightforward purpose.
Evolution.
@@
The evolution step involves the continuous improvement and adaptation of the API based on outcomes and new requirements in terms of diversity and complexity. An LLM is instructed to synthesize new APIs according to a sampled subtree of the API context tree and an API example. The generated definitions of new APIs are required to be clear and thorough. We then apply a set of diversity indicators, e.g., adding new functionalities or parameters, including additional constraints, mutating parameter type, and updating returned results, to diversify the generated APIs. We maintain an API example buffer containing various API examples. Iteratively, we sample an example from the buffer, adapt it to the current subtree of functionalities, and generate the next generation of the APIs.
@@
The proposed TSS module facilitates the efficient generation of a diverse and complex set of API documentation, with nested types including lists of lists or lists of dictionaries. Through this process, we have curated a collection of 26,507 distinct APIs.
Figure 2: The detailed process of the Tool Self-Evolution Synthesis (TSS). The left part shows a subtree of the API context tree as an example.
2.2 Multi-Agent Interactive Dialog Generation
Real-world tasks involving function calls often encompass complex intents and varied requirements. To solve these real-world tasks, tool-augmented LLMs are expected to accurately identify when, what, and how many function calls are needed. To better represent real-world scenarios, we propose a multi-agent interactive (MAI) dialog generation module to synthesize function-calling dialogs. These dialogs cover a wide range of types, including simple function calls, parallel function calls, dependent functions, and non-tool-use dialogs. The MAI module is designed to ensure accuracy, complexity, and diversity throughout the generation process.
2.2.1 Dialog Generation Overview
The middle part of Figure 1 illustrates our dialog generation process. Initially, we sample one or more API candidates from our curated API pool, ensuring that the selected APIs share the same domain. This helps maintain topic consistency within a single dialog. Additionally, APIs within the same domain are more likely to have similar functions. Dialogs based on similar APIs can enhance the model’s ability to make nuanced distinctions between different APIs.
@@ 2.2.1 Dialog Generation Overview
We then generate dialogs through the interplay among three different agents (user, assistant, and tool), each simulated by an LLM. The user agent mainly makes requests or provides additional information, powered by multi-mode prompting and similarity-guided complication to enhance diversity and complexity. The assistant agent addresses the user’s queries equipped with the given APIs. The action space of the assistant agent includes: calling the APIs, requesting further information, summarizing the tool feedback, and providing non-tool-use answers. Each action is determined through a formalized thinking process and validated by self-consistency to ensure accuracy. The tool agent acts as the API executor, processing tool descriptions and input parameters provided by the assistant, and outputs the potential execution results. Finally, after the dialogs are generated, they can be converted to diverse formats with our designed tool format generalization. Algorithm 1 outlines the overall dialog generation process.
In the following sections, we will elaborate on the MAI module and explain how accuracy, diversity, and complexity are ensured during the generation process.
Algorithm 1 MAI Dialog Generation
1:Initialization: Sampled API list $A$, Dialog $D_{0}=[]$, Target Turn Length $N_{t}$
2:Definition: User Agent $\theta_{u}$ and its output $C_{u}$, Assistant Agent $\theta_{a}$ and its output $C_{a}$, Tool Agent $\theta_{t}$ and its output $C_{t}$
3:for $t=1,2,\ldots,N_{t}$ do
4:     Sample user template $p$ and user style $s$
5:     $C_{u}=\theta_{u}(D_{t-1},A,p,s)$
6:     $C_{a}^{1},C_{a}^{2},\ldots,C_{a}^{N}=\theta_{a}(C_{u},D_{t-1},A)$triangleright Generate response N times with formalized thinking.
7:     if $C_{a}^{1}\neq C_{a}^{2}\neq\ldots\neq C_{a}^{N}$ then $\triangleright$ Only check consistency of tool callings.
8:         Continue or Add Loss Mask $\triangleright$ Discard the turn or add a loss mask to the turn.
9:     else
10:         $C_{a}=\text{MajorVote}(C_{a}^{1},C_{a}^{2},\ldots,C_{a}^{N})$
11:     end if
12:     $D_{t}=D_{t-1}+[C_{u},C_{a}]$
13:     while Tool calling in $C_{a}$ do $\triangleright$ Dependent functions need several sequential tool callings.
14:         $C_{t}=\theta_{t}(C_{a},A)$
15:         $C_{a}=\theta_{a}(C_{t},D_{t},A)$
16:         $D_{t}=D_{t}+[C_{t},C_{a}]$
17:     end while
18:end $\triangleright$ The generated dialog $D_{N_{t}}$ can be further transformed to diverse formats.
2.2.2 Generation Ensuring Accuracy, Diversity and Complexity
The MAI module generates function-calling dialogs using multiple strategies to ensure accuracy, diversity, and complexity. We introduce these strategies in sequence below.
Multi-Mode Prompting.
MAI generation begins by determining the target dialog mode for each sample. Our multi-mode prompting allows us to create a wide range of dialog types by varying the instructions given to the user agent. We generate both single-turn and multi-turn dialogs, guided by the specified turn length in the instructions.
To further diversify the dialog types, we adjust the number of tool calls required and the relationships among these calls during prompting, resulting in single, parallel, and dependent function calls. Additionally, we produce non-tool-use dialogs, categorized into two scenarios: when no suitable tools are available and when there is insufficient information to call the tools.
This variety of dialogs is essential for developing a comprehensive skill set in tool usage for LLMs.
Similarity-Guided Complication.
Given guidance from the instructions, the user agent is able to generate an appropriate query requiring demanded function types. Apart from that, we also consider complexity in terms of linguistic level, which can be identified based on a similarity metric (please refer to Section 4.2 for details). To capture queries of varying complexity, we employ a range of templates to prompt the user agent with different linguistic styles. The query’s complexity can be calculated after the dialog concludes, and then used as a criterion for data selection.
Formalized Thinking.
After the user agent generates the query, the assistant agent must determine the appropriate response action. Research has shown that encouraging LLMs to think before acting – such as through chain-of-thought prompting [21] – enhances their reasoning abilities. We employ a similar strategy to improve the accuracy of tool-calling decision-making for our assistant agent. Unlike traditional chain-of-thought prompts, which typically encourage general reflective thinking (e.g., by merely adding "Let’s think step by step"), we implement a specialized and structured thinking process tailored specifically for function calls. This process involves: 1) evaluating the user query, 2) assessing the tool requirements, and 3) ensuring all required parameters are provided.
Formalized Thinking is utilized only during data generation and is excluded from training data formation to maintain conciseness.
Self-Consistency.
To further ensure accuracy, we implement a self-consistency mechanism when deciding actions. Each response from the assistant agent is generated multiple times. A response is adopted only if the action decisions (including whether to make tool calls and the parameter values to be filled) are consistent across at least two instances. If consistency is achieved, we apply majority voting to select the final response. If consistency is not met, the current turn of the dialog is either discarded or a loss mask is applied to prevent the model from learning potentially erroneous content.
@@ 2.2.2 Generation Ensuring Accuracy, Diversity and Complexity
If the action determined by the assistant agent involves tool calls, the tool agent will provide simulated results, which the assistant agent then summarizes and presents to the user (for queries requiring dependent function calls, additional interactions between the assistant and tool agents are necessary). The generation process will continue with the user agent querying again or responding to the assistant’s question until the target turn length is reached.
Tool Format Generalization.
Format generalization can be applied to the generated dialogs at the end to increase format diversity. All generated dialogs are initially stored in JSON format, which is easy to process and validate (details will be shown in next section). To support a flexible and user-specific format for tool definitions and function calls, we further convert our generated dialogs and API lists into various formats (e.g., JSON, YAML, XML, Markdown). We observe that the models’ instruction-following capabilities improve as the format diversity increases.
3 Data Verification

### v2

@@ Abstract
Function calling significantly extends the application boundary of large language models (LLMs), where high-quality and diverse training data is critical for unlocking this capability. However, collecting and annotating real function-calling data is challenging, while synthetic data from existing pipelines often lack coverage and accuracy. In this paper, we present ToolACE, an automatic agentic pipeline designed to generate accurate, complex, and diverse tool-learning data, specifically tailored to the capabilities of LLMs. ToolACE leverages a novel self-evolution synthesis process to curate a comprehensive API pool of 26,507 diverse APIs. Dialogs are further generated through the interplay among multiple agents, under the guidance of a complexity evaluator. To ensure data accuracy, we implement a dual-layer verification system combining rule-based and model-based checks. We demonstrate that models trained on our synthesized data—even with only 8B parameters—achieve state-of-the-art performance, comparable to the latest GPT-4 models. Our model and a subset of the data are publicly available at https://huggingface.co/Team-ACE.
@@ 1 Introduction
Equipping Large Language Models (LLMs) with external tools has significantly enhanced the capability of AI Agents to solve complex real-world tasks Huang et al. (2024); Qin et al. (2023); Qu et al. (2024). The integration of function calling enables LLMs to access up-to-date information, perform delicate computations, and utilize third-party services, thereby unlocking a wide range of potential applications across various fields, e.g., workflow automation Zhong et al. (2023), financial reporting Theuma & Shareghi (2024), and travel planning Hao et al. (2024).
Function calls in real-world applications are often diverse and complex, driven by the varied functionalities of APIs [^1] In this paper, APIs, tools, functions, and plugins are used interchangeably. and the broad range of tasks Qin et al. (2023). APIs often undergo rapid updates to meet diverse user needs, necessitating models capable of robust zero-shot generalization. Additionally, users’ requirements can be complex or ambiguous, leading to scenarios where multiple tools are employed in a parallel or dependent manner, or require multi-turn interactions. This highlights the importance of managing intricate instructions and accommodating various function-calling scenarios.
Despite these challenges, current tool-augmented LLMs primarily focus on simple function-calling tasks with limited diversity and complexity Qu et al. (2024). They mainly rely on existing public APIs for task construction, which restricts their zero-shot capabilities and applicability to single-turn queries, neglecting more complex scenarios such as dependent or multi-turn interactions Qin et al. (2023); Tang et al. (2023); Liu et al. (2024). Table 1 provides an overview of the data statistics used in these representative tool-augmented LLMs. Moreover, executions of function calls demand precise API selection and parameter configuration, which are highly dependent on the quality and accuracy of underlying data. As data becomes increasingly diverse and complex, generating accurate samples with simple pipelines introduced by the existing work becomes significantly more challenging.
@@
Gorilla Patil et al. (2023) 1645 3
ToolAlpaca Tang et al. (2023) 3938 50
ToolLLM Qin et al. (2023) 16464 49
Functionary Meetkai (2024) n/a n/a
xLAM Liu et al. (2024) 3673 21
Granite Abdelaziz et al. (2024) n/a n/a
@@
In this paper, we present ToolACE, a systematic tool-learning pipeline that automatically synthesizes accurate, diverse, and complex function calling data, with the awareness of the model’s capability.
Evolutionary Diversity. Exposing LLMs to a broad range of function-calling scenarios enhances their proficiency and zero-shot tool usage Zhang et al. (2024). Instead of relying on public APIs, ToolACE introduces a Tool Self-Evolution Synthesis (TSS) method. TSS uses a speciation-adaptation-evolution process to generate tools across multiple domains with diverse data types and constraints. Starting with pretraining data to ensure comprehensive coverage, this iterative process of self-evolution and continual updates expands the diversity of the API pool, enabling more sophisticated data generation.
Self-Guided Complexity. Instruction-following data should possess sufficient complexity to foster function-calling skills. LLMs learn more effectively when the complexity of the data slightly exceeds their current capability Du et al. (2023). To address this, we propose a self-guided dialog generation process (SDG), where the LLM serves as an evaluator to regulate complexity. Four types of function-calling data are generated with multi-agent interactions, following a self-guided complication strategy.
Refined Accuracy. Data accuracy is fundamental to the effectiveness of tool-augmented LLMs. ToolACE employs a dual-layer verification (DLV) system, integrating both rule-based and model-based checkers, to guarantee the executability and consistency of the synthesized data.
Equipped with data accuracy, complexity, and diversity, ToolACE aims to enhance the function-calling capability of LLMs with strong generalization. Our contributions are outlined as follows:
  • We propose a novel automated data pipeline for function calls, ToolACE, which comprises a tool self-evolution synthesis module, a self-guided dialog generation module, and a dual-layer verification module. To our knowledge, this is the first work to highlight the benefits of synthesizing diverse APIs to improve the generalization of function calls.
  • We develop a self-guided complication strategy to generate various types of function-calling dialogs with appropriate complexity. The given LLM is utilized as the complexity evaluator to guide the complexity level of the generated data. The quality of the generated data is ensured through a dual-layer verification process, which combines both rule checkers and model checkers.
  • We conduct experiments on two widely adopted benchmarks: BFCL Yan et al. (2024) and APIBank Li et al. (2023). With only 8B parameters, ToolACE significantly outperforms existing open-source LLMs and is competitive with the latest GPT-4 models.
@@ 2 Data Generation Pipeline
Effective use of synthetic data significantly enhances the capabilities of large language models (LLMs) Mitra et al. (2024). Hence, in ToolACE, we propose an automated agentic framework for tool learning to generate high-quality, diverse, and complex data, guided by the capability of the given LLM to be tuned, as illustrated in Figure 1. The proposed framework deploys various agents to recursively synthesize diverse APIs, collaboratively construct dialogs with appropriate complexity, and rigorously reflect on data quality. The following sections present our Tool Self-evolution Synthesis (TSS) module, Self-Guided Dialog Generation (SDG) module, and Dual-Layer Validation Process (DLV).
@@ 2.1 Tool Self-evolution Synthesis
The variety of APIs significantly underpins the diversity of the function-calling data. As shown in Table 1, ToolACE has established a comprehensive API pool that surpasses other representative tool-augmented LLMs in both quantity and domain coverage, incorporating both real and synthesized APIs. Beyond collecting real API data, we developed a Tool Self-Evolution Synthesis (TSS) module that synthesizes API definitions with various data types and constraints, which encompasses three major steps: 1) Speciation, 2) Adaptation, and 3) Evolution.
@@
Figure 1: The overall framework of ToolACE, which mainly consists of Tool Self-evolution Synthesis (TSS), Self-Guided Dialog Generation (SDG), and Dual-Layer Validation Process (DLV).
@@
Speciation. APIs with extensive domain coverage enable tool-augmented LLMs to learn a wider array of use cases from various applications and industries, thereby significantly enhancing their generalization ability. In the speciation step, we propose to create a hierarchical API context tree to guide the synthesis process with possible API domains and functionalities.
@@
We observe that the pretraining data for LLMs encompasses one of the most diverse sources of human corpus, providing a solid foundation for extracting various API domains and use cases. Starting with API-related raw documents from the pretraining data (e.g., technical manuals, API documentation, product specifications, user guides, and tutorials), we prompt an agent powered by a frontier LLM to extract an API domain along with all possible API functionalities or use cases from each document. Children nodes of the context tree are recursively generated at each step, with each node denoting a possible API functionality (e.g., get the weather forecast, get the stock price, send an email). Figure 9 in the Appendix A showcases the subtree under the entertainment domain as an example.
@@
Adaptation. In the adaption step, we specify the domain and diversity level of each API. We sample a subtree and obtain unique functionalities from the API context tree for each individual API, so that different APIs possess distinct functionalities. For example, some APIs may cover more nodes, thereby acquiring more domain-specific and detailed capabilities. Whereas some APIs may only include a single node from the context tree, focusing on an easy, straightforward purpose.
@@
Evolution. The evolution step involves the continuous improvement and adaptation of the API based on outcomes and new requirements. An LLM is instructed to synthesize new APIs according to a sampled subtree of the API context tree and an API example. The generated definitions of new APIs are required to be clear and thorough. We then apply a set of diversity indicators, e.g., adding new functionalities or parameters, including additional constraints, mutating parameter type, and updating returned results, to diversify the generated APIs. We maintain an API example buffer containing various API examples. Iteratively, we sample an example from the buffer, adapt it to the current subtree of functionalities, and generate the next generation of the APIs.
@@
The proposed TSS module facilitates the efficient generation of a diverse set of API documentation, with nested types including lists of lists or lists of dictionaries.
2.2 Self-Guided Dialog Generation
The effectiveness of function-calling data is closely tied to the capabilities of the LLM. For different LLMs, the knowledge and abilities they have learned during the pretraining phase are different, thereby the function-calling data they require should also differ Du et al. (2023). For instance, an LLM with 0.5B parameters may struggle to comprehend extremely complex data with long dependencies between APIs. In contrast, a well-trained 70B LLM can easily handle straightforward queries with clear intentions and simple APIs. In both cases, the data is unproductive for the given LLM, highlighting the importance of tailoring data generation to align with the model’s capabilities.
Hence, to ensure the generated dialogs indeed fill the ability gap for the given LLM, we propose a self-guided dialog generation (SDG) module to synthesize the function-calling dialogs, as shown in the middle part of Figure 1. SDG consists of a complexity evaluator and a multi-agent generator. Various types of function-calling dialogs are generated via the interaction of multiple agents. The LLM to be tuned serves as the evaluator, assessing the complexity of the generated data. Data that is deemed too simple or too complex is dynamically adjusted under the guidance of the evaluator.
2.2.1 Multi-agent Dialog Generation
We propose a multi-agent framework to generate the four types of function-calling dialogs: single function calls, parallel function calls, dependent function calls, and non-tool-use dialogs.
@@ 2.2.1 Multi-agent Dialog Generation
The data generator includes three agents—user, assistant, and tool—each simulated by an LLM. One or more API candidates are sampled from our curated API pool and present the sampled APIs to the agents. Dialogs are then generated through role-playing among the three agents, each agent is provided with a necessary role assignment and detailed task description to continue the conversation. The user agent mainly makes requests or provides additional information to the assistant, with a self-guided complication process to adjust the dialog complexity. The assistant agent addresses the user’s queries equipped with the given APIs. The action space of the assistant agent includes: calling the APIs, requesting further information, summarizing the tool feedback, and providing non-tool-use answers. To ensure data quality, each assistant action is generated multiple times, and only responses with consistent decisions across multiple instances are adopted. A specialized and structured thinking process specifically designed for function calls is also applied to enhance the assistant’s tool-calling decisions. The tool agent acts as the API executor, processing tool descriptions and input parameters provided by the assistant, and outputs the potential execution results.
@@
For each function-calling dialog, the user agent initiates a request related to the given sampled APIs. The assistant agent reviews the request and decides whether to call an API or ask for additional information. If tool calls are required, the tool agent will provide simulated results, and the assistant agent will summarize the results and present the user. The generation process continues with the user agent querying again or responding to the assistant’s question until the target turn length is reached.
2.2.2 Data Complexity Evaluation
Different LLMs exhibit varying knowledge and capabilities, which necessitates the use of different data to optimize tool usage performance. However, much of the existing research overlooks the correlation between the model capability and the training data, leading to suboptimal data efficiency.
[…]
$H_{\mathcal{M}}(x,y)=-\frac{1}{n_{y}}\sum_{i=1}^{n_{y}}\log p(t_{i}|x,t_{1},\dots,t_{i-1})\,,$ (1)
[…]
Our findings suggest that the loss of a data sample is generally positively correlated with (1) the number of candidate APIs available for selection, (2) the number of APIs utilized, and (3) the dissimilarity between the user query and the API descriptions, as demonstrated in Figure 2. Intuitively, as the number of candidate APIs increases, selecting the correct one becomes more difficult. Similarly, the use of a higher number of APIs reflects greater query complexity, while larger discrepancies between the user query and the API descriptions demand more sophisticated reasoning to identify the correct function. These validate the use of loss as a measure of data complexity in function calling.
To establish an appropriate complexity range for the given LLM $\mathcal{M}$, we create a small, prior data set that spans various levels of complexity. A data sample that is correctly generated by $\mathcal{M}$ indicates that the model has already mastered the corresponding tool usage case, and thus this sample is unnecessary for further fine-tuning. The associated loss serves as a reference lower bound for data complexity. Conversely, if the loss of a data sample remains high after fine-tuning, it may indicate that the sample is too complex for the model to learn, and this loss serves as a reference upper bound.
Our evaluator provides the suitable complexity range, along with the loss of the given data sample, as the guidance information for the multi-agent generator in generating the training data.
(a) Number of candidate APIs
(b) Number of utilized APIs
(c) Dissimilarity
Figure 2: Relationships between loss and (1) the number of candidate APIs available for selection, (2) the number of APIs utilized, and (3) the dissimilarity between the user query and the API descriptions.
2.2.3 Self-Guided Complication
After obtaining the complexity of the current data from the evaluator, the user agent’s instructions are dynamically adjusted. If the data sample is too simple for the LLM, the user agent is instructed to generate a more complex query–one that either requires additional APIs or diverges further from the API description to increase complexity. Conversely, if the data sample exceeds the LLM’s capacity, the user agent is prompted to produce a simpler query. In this way, the data generation process is continually adapted to better match the model’s performance level.
2.3 Dual-Layer Data Verification